Three pipes can be used to fill a tank. It is filled in 10 hours using only the first pipe,in 12 hours using only the second pipe and in 15 hours using only the third one. A pressure problem cuts the flow capacity of each pipe in half. It is then decided to use three pipes together. How long will it take to fill a tank?
(A)4 hrs (B) 5 hrs (C) 8 hrs (D) 12 hrs (E) None of these

Answer by stanbon(75887)

Three pipes can be used to fill a tank. It is filled in 10 hours using only the first pipe,in 12 hours using only the second pipe and in 15 hours using only the third one. A pressure problem cuts the flow capacity of each pipe in half. It is then decided to use three pipes together. How long will it take to fill a tank?
---
Since pressure is halfed the times for each is doubled:
-----
1st pipe DATA:
time = 20 hr/job ; rate = 1/20 job/hr
------------------------
2nd pipe DATA:
time = 24 hr/job ; rate = 1/24 job/hr
------------------------
3rd pipe DATA:
time = 30 hr/job ; rate = 1/30 job/hr
-------------------------
Together DATA:
time = x hr/job ; rate = 1/x job/hr
============================================
Equation:
rate + rate + rate = together rate
1/20 + 1/24 + 1/30 = 1/x
(24*30x) + (20*30x) + (20*24x) = 20*24*30
x = 8 hrs.
=========================================
Cheers,
Stan H.




(A)4 hrs (B) 5 hrs (C) 8 hrs (D) 12 hrs (E) None of these